Friday, August 31, 2012

Night Log 2012-08-30

SA: Tim
SO: Thea
Others: luke h


first half of the night was breezy with come and go cirrus. winds picked up in the 2nd half and closed us down eventually shortly after 0300. only successful science data was FT spectroscopy of V2301 Oph for DoD. attempted a couple of FP targets, but bright moon and poor IQ scuttled them. got full sets of FP calibration data for the entire H-alpha region.

mirror alignment was once again a struggle and sucked up significant time during both attempts tonight.

the timdimm is working now, but only the dimm portion. still need to sort out a computer to run the mass.

No comments:

Post a Comment